10. David Corporation uses the weighted-average method in its process costing system. The first processing department, the Welding Department, started the month with 20,000 units in its beginning work in process inventory that were 80% complete with respect to conversion costs. The conversion cost in this beginning work in process inventory was $123,200. An additional 65,000 units were started into production during the month. There were 19,000 units in the ending work in process inventory of the Welding Department that were 10% complete with respect to conversion costs. A total of $389,250 in conversion costs were incurred in the department during the month.
What would be the cost per equivalent unit for conversion costs for the month? (Round off to three decimal places.)

_________________________________________________________________________________________________

11. Jimmy Corporation uses the weighted-average method in its process costing system. The ending work in process inventory consists of 9,000 units. The ending work in process inventory is 100% complete with respect to materials and 70% complete with respect to labor and overhead. If the cost per equivalent unit for the period is $3.75 for material and $1.25 for labor and overhead, what is the balance of the ending work in process inventory account?
Answer #1

10) Unit transferred out = 20000+65000-19000 = 66000

Equivalent unit of conversion = 66000+(19000*10%) = 67900

Cost per equivalent unit of conversion = (123200+389250)/67900 = 7.547

11) Cost of ending work in process = (9000*3.75+9000*70%*1.25) = 41625